# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 9.5 Conditioning Technique
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{t+1}{})^{2},$ (61)
@@
    Next, we derive (36b) for $r=t$ and $10\leq s\leq t-$ 1. Considering (33), we can write
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\Upsilon^{i}_{s+1}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\Upsilon^{i}_{s+1}{}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\Upsilon^{i}_{s+1}{}\right).$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\Upsilon^{i}_{t+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\Upsilon^{i}_{s+1}{}$ (63)
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\sum_{r=0}^{t-1}\gamma^{r}_{t}\left(\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}\right)$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\left(\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{r=0}^{t-1}\gamma^{r}_{t}U^{i}_{r}{}U^{i}_{s}{}\right)$
@@
[…]
    The derivations for (36c) and (36d) follow a similar procedure, which we omit here for brevity. We then apply a similar approach to obtain (36e). Specifically, fixing $10\leq r\leq t-$ 1 and setting $s=t$, we can write the following using (33) and (63):
@@
    Likewise, we can show the result for the case that $r=t$ and $10\leq s\leq t-$ 1:
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{},$
[…]
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]}.$
@@
    $\Psi^{i}(N)$ $:=\psi\big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big{)},$
@@
    $\lim_{N\rightarrow\infty}\tilde{\nu}_{(t+1)N}^{i}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[({M}^{i}+{M}_{t}^{i})g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big{)}\right]$ (67)
@@
    $(\hat{\rho}_{tN})^{2}$ $:=\mathrm{Var}\left[[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}]^{i}\Big{|}\vec{U}_{t}{}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}+\frac{\sigma_{t}^{2}}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2},$ (68)
@@
    $\frac{1}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}=\left<\vec{U}^{\perp}_{t}{},\vec{U}^{\perp}_{t}{}\right>$ =$\vec{U}_{t}{},\vec{U}_{t}{}\right>-\left<\vec{U}^{\parallel}_{t}{},\vec{U}^{\parallel}_{t}{}\right>=\frac{1}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2}-\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}.$ (71)
@@
    $\quad\quad\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}=\lim_{N\rightarrow\infty}\sum_{r,s=0}^{t-1}\gamma^{r}_{t}\gamma^{s}_{t}\left<\vec{U}_{r}{},\vec{U}_{s}{}\right><\infty.$ (73)
@@
    where $c$ is a constant independent of $N$ and we used the inequality $(v_{1}+v_{2})^{l}\leq 2^{l-1}(v_{1}^{l}+v_{2}^{l}),\;v_{1},v_{2}\geq 0$. Note that in (74), given ${\cal G}_{t}$, the term $\bm{R}_{t}\vec{\gamma}_{t}$ is deterministic and bounded by Corollary 9.13. Then, fixing $10<\kappa<$ 1 and using the fact that $\psi\in{{\cal C}{\cal P}}(k)$ and so $|\psi(\vec{\omega})|\leq c(1+{\left\|\vec{\omega}\right\|}^{k})$, by Assumption 9.4, we get,
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\;\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{{\bf A},{\bf B}_{t}}\Big{[}\psi$ $\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\hat{\nu}_{(t+1)N}^{i}{}\Big{)}\Big{]}.$
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{Z}\Big{[}\psi\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big{)}\Big{]},$
@@
    $:=\mathbb{E}_{Z}\Big{[}\psi$ $\Big{(}Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big{)}\Big{]}.$
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg{[}\widehat{\psi}\bigg{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\bigg{)}\bigg{]}$
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg{[}\psi\bigg{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},\tilde{\nu}_{t+1}^{\mathcal{S}}{}\bigg{)}\bigg{]},$
@@
    Whenever $\psi\in{{\cal C}{\cal P}}(k)$, by Assumption 9.4, we get that $\widetilde{\psi}\in{{\cal C}{\cal P}}(k)$, and applying the result in Eq. (78) for the function $\widetilde{\psi}$ yields the desired result. $\Box$

### v2

@@ 10.5 Conditioning Technique
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{t+1}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{t+1}{})^{2},$ (66)
@@
    Next, we derive (41b) for $r=t$ and $0\leq s\leq t-1$. Considering (38), we can write
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\Upsilon^{i}_{s+1}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\Upsilon^{i}_{s+1}{}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\Upsilon^{i}_{s+1}{}\right).$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{s+1}{}\Upsilon^{i}_{t+1}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}\right)\Upsilon^{i}_{s+1}{}$ (68)
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\sum_{r=0}^{t-1}\gamma^{r}_{t}\left(\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}\right)$
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\hskip-2.84526pt\left(\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{r=0}^{t-1}\gamma^{r}_{t}U^{i}_{r}{}U^{i}_{s}{}\right)$
@@
    where in the last line, we used the fact that $\left<\vec{U}_{t}{},\vec{U}_{s}{}\right>=\left<\vec{U}^{\parallel}_{t}{},\vec{U}_{s}{}\right$> as $\vec{U}^{\perp}_{t}{}\perp\vec{U}_{s}{}$.
    The derivations for (41c) and (41d) follow a similar procedure, which we omit here for brevity. We then apply a similar approach to obtain (41e). Specifically, fixing $0\leq r\leq t-1$ and setting $s=t$, we can write the following using (38) and (68):
@@
    Likewise, we can show the result for the case that $r=t$ and $0\leq s\leq t-1$:
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big)\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)$
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{},$
    where in the last line, we used $\left<\vec{U}_{t}{},\vec{U}_{s}{}\right>=\left<\vec{U}^{\parallel}_{t}{},\vec{U}_{s}{}\right$>. The desired result follows by aggregating (69)-(70).
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[\psi\big(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big)\Big].$
@@
    $\Psi^{i}(N)$ $=\psi\big(Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\Upsilon^{i}_{t+1}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{(t+1)N}^{i}{}\big),$
@@
    $\lim_{N\rightarrow\infty}\tilde{\nu}_{(t+1)N}^{i}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[({M}^{i}+{M}_{t}^{i})g_{t}\big(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z+H_{t-1}{},\vec{W},{\vec{X}}\big)\right]$ (72)
@@
    $(\hat{\rho}_{tN})^{2}$ $=\mathrm{Var}\left[[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+{\bf B}_{t}\vec{U}_{t}{}]^{i}\Big|\vec{U}_{t}{}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}+\frac{\sigma_{t}^{2}}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2},$ (73)
@@
    $\frac{1}{N}\sum_{j=1}^{N}\left(U^{\perp,j}_{t}{}\right)^{2}=\left<\vec{U}^{\perp}_{t}{},\vec{U}^{\perp}_{t}{}\right$> $=\left<\vec{U}_{t}{},\vec{U}_{t}{}\right>-\left<\vec{U}^{\parallel}_{t}{},\vec{U}^{\parallel}_{t}{}\right>=\frac{1}{N}\sum_{j=1}^{N}\left(U^{j}_{t}{}\right)^{2}-\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}.$ (76)
@@
    $\quad\quad\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\left(U^{\parallel,j}_{t}{}\right)^{2}=\lim_{N\rightarrow\infty}\sum_{r,s=0}^{t-1}\gamma^{r}_{t}\gamma^{s}_{t}\left<\vec{U}_{r}{},\vec{U}_{s}{}\right$$\infty$. (78)
@@
    where $c$ is a constant independent of $N$ and we used the inequality $(v_{1}+v_{2})^{l}\leq 2^{l-1}(v_{1}^{l}+v_{2}^{l}),\;v_{1},v_{2}\geq 0$. Note that in (79), given ${\cal G}_{t}$, the term $\bm{R}_{t}\vec{\gamma}_{t}$ is deterministic and bounded by Corollary 10.14. Then, fixing $0<\kappa<1$ and using the fact that $\psi\in{{\cal C}{\cal P}}(k)$ and so $|\psi(\vec{\omega})|\leq c(1+{\left\|\vec{\omega}\right\|}^{k})$, by Assumption 10.4, we get,
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\;\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{{\bf A},{\bf B}_{t}}\Big[\psi$ $\Big(Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\left[\widetilde{{\bf A}}\vec{U}^{\perp}_{t}{}+\bm{R}_{t}\vec{\gamma}_{t}+{\bf B}_{t}\vec{U}_{t}{}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\hat{\nu}_{(t+1)N}^{i}{}\Big)\Big].$
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{1}{\left|{\mathcal{S}}\right|}\sum_{i\in\mathcal{S}}\mathbb{E}_{Z}\Big[\psi\Big(Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big)\Big],$
@@
    $=\mathbb{E}_{Z}\Big[\psi$ $\Big(Y^{i}_{0}{},Y^{i}_{1}{},\ldots,Y^{i}_{t}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big),\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{t+1}^{i}{}\Big)\Big].$
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg[\widehat{\psi}\bigg(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\bigg)\bigg]$
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\;\mathbb{E}\bigg[\psi\bigg(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\hat{\rho}_{t}Z+\sum_{s=0}^{t-1}\gamma^{t}_{s}\bar{\rho}_{s+1}{}Z^{\prime}_{s+1},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},\tilde{\nu}_{t+1}^{\mathcal{S}}{}\bigg)\bigg],$
@@
    Whenever $\psi\in{{\cal C}{\cal P}}(k)$, by Assumption 8, we get that $\widetilde{\psi}\in{{\cal C}{\cal P}}(k)$, and applying the result in Eq. (83) for the function $\widetilde{\psi}$ yields the desired result. $\Box$